### v1

@@ 3 Subspace method based on neural networks for eigenvalue problems
$\begin{cases}Lu=\lambda u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega,\end{cases}$
where the differential operator $L:=-\nabla\cdot(\alpha\nabla)+\beta$ and $\alpha=(\alpha_{ij})_{ij}$ is symmetric positive definite with $\alpha_{ij}\in W^{1,\infty}(\Omega)(i,j=1,\cdots,d)$, and $0\leq\beta\in L^{\infty}(\Omega).$
The associate weak form of the eigenvalue problem above is as follows: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\Omega)$ such that $b(u,u)=1$ and
$a(u,v)=\lambda b(u,v),\quad\forall v\in H_{0}^{1}(\Omega),$ (5)
which is nothing but (1) with $a(u,v)=(\alpha\nabla u,\nabla v)+(\beta u,v)$ and $b(u,v)=(u,v)$. Here, $a(\cdot,\cdot),b(\cdot,\cdot)$ are two positive definite symmetric bilinear forms on $H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)$.
@@
$\|v\|_{1}^{2}-C\|v\|_{0}^{2}\lesssim a(v,v),\quad\forall v\in H_{0}^{1}(\Omega)$
@@
3.1 A general framework of SNN for eigenvalue problems
In this subsection, we extend the subspace method based on neural networks for PDEs, which is proposed in [49] and introduced in Section 2, to eigenvalue problems.
A neural networks architecture similar to the one introduced in the previous section is adopted. The only difference is that the neural network in Section 2 is used to approximate the solution to the boundary value problem, here the neural network is used to approximate the eigenfunctions corresponding to the first $k$ eigenvalues of some partial differential operator. Therefore, the dimension of the output layer is $k$. Thus, the output of the neural network can be presented as follows,
$\bm{u}(x;\theta,\bm{w})=\begin{pmatrix}u_{1}(x;\theta,w_{1})\\ \cdots\\ u_{k}(x;\theta,w_{k})\end{pmatrix}=\begin{pmatrix}\varphi\cdot w_{1}\\ \cdots\\ \varphi\cdot w_{k}\end{pmatrix}=\begin{pmatrix}\sum\limits_{j=1}^{M}w_{1}^{(j)}\varphi_{j}(x,\theta)\\ \cdots\\ \sum\limits_{j=1}^{M}w_{k}^{(j)}\varphi_{j}(x,\theta)\end{pmatrix}$
where $\bm{w}=(w_{1},\cdots,w_{k})$, $w_{s}=(w_{s}^{(1)},\cdots,w_{s}^{(M)})^{T}$ denote the coefficients corresponding to the $s$-th output, and $u_{s}(\bm{x};\theta,w_{s})$ denotes the $s$-th output, $s=1,\cdots,k$.
Following a framework similar to Algorithm 1, we initialize $\theta$ and $w_{1},\cdots,w_{k}$ firstly. Then, to handle the boundary condition, appropriate modifications to the basis functions are made according to the specific problem requirements. Next, we fix the coefficients $w_{1},\cdots,w_{k}$ and update parameters $\theta$ by minimizing some loss function, from which we obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$. Finally, we calculate the Galerkin projection of the original eigenvalue problem onto the subspace spanned by basis functions $\varphi_{1},\cdots,\varphi_{M}$ and obtain an approximate solution to the eigenvalue problem (5). The above process is summarized as Algorithm 2.
@@ 3.1 A general framework of SNN for eigenvalue problems
Algorithm 2 General framework of SNN for eigenvalue problems
  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w_{1}$$\cdots$,$w_{k}$, then modify the neural network basis functions according to the boundary conditions.
  2. Update parameters $\theta$ by minimizing a loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
  3. Calculate the Galerkin projection of the eigenvalue problem (5) onto the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ and obtain an approximate solution.
In the following subsections, we will follow the process outlined in the Algorithm 2 to provide a step-by-step explanation.
3.2 Initializing and modifying basis functions of the subspace
3.2.1 Initializing the basis functions
@@
First, we introduce the initialization of the neural network parameters $\theta$ and the coefficients $w_{1},\cdots,w_{k}$. There are various ways to initialize the parameters $\theta$ for initializing the neural network. In this paper, we adopt the default initialization strategy provided by PyTorch for neural network parameters $\theta$. The coefficients $w_{1},\cdots,w_{k}$ are given and will remain fixed during the training process. While there are various methods for assigning the coefficients $w_{1},\cdots,w_{k}$, numerical experiments provided in [49] indicate that the different choices of $w_{1},\cdots,w_{k}$ have no obvious impact on the results. Therefore, here, we do not pay much attention on the choice of $w_{1},\cdots,w_{k}$. We just give some simple choices. That is, for a single output, we choose $w=(1,\cdots,1)^{T}\in\mathbb{R}^{M}$. For multiple outputs, a random approach is chosen, that is, $w_{1},\cdots,w_{k}\in\mathbb{R}^{M}$ are assigned random values generated from a uniform distribution defined over the interval $[-1,1]$.
3.2.2 Modifying the basis functions
@@
For a partial differential equation, some boundary conditions are usually required. Since the basis functions constructed by some neural networks are usually not local, it is not so easy to deal with the boundary condition by using the same way as traditional methods. Here, we choose to modify the basis functions to make them satisfy the boundary condition. For example, for the case of zero boundary condition with $\Omega$ being bounded, we can modify the basis function by the following way. We multiply $\varphi_{i}$ by a function $f$ such that
$\bar{\varphi}_{i}=\varphi_{i}f,\ \bar{\varphi}_{i}(\bm{x})=0,\ \forall\bm{x}\in\partial\Omega,\ \forall i=1,\cdots,M.$
@@
If $\Omega$ is unbounded, it is necessary to modify the basis functions to ensure that the modified functions are $L^{2}$ integrable. We can achieve this by multiplying $\varphi_{i}$ by a function $f$ such that
[…]
Since the Hermite-Gauss quadrature scheme are used to numerically approximate the integrals on unbounded domain in this paper, we can choose $f(\bm{x})=\exp(-\frac{1}{2}\bm{x}^{T}\bm{x})$.
For simplicity, in the remaining parts of this paper, $\varphi_{i}$ will refer to the modified basis function $\bar{\varphi}_{i}$ as described above.
3.3 Updating the basis functions
In this subsection, we will introduce a specific approach to update basis functions by minimizing some loss function $\mathcal{L}$.
Denote $\bm{w}=(w_{1},\cdots,w_{k})$. For given coefficients $\bm{w}$, parameters $\theta$ can be updated by solving the following minimization problem:
$\min_{\theta\in\Theta}\mathcal{L}(\theta,\bm{w}).$ (6)
The gradient descent method is adopted to solve minimization problem (6). That is, we update the parameter by
$\theta^{(l+1)}=\theta^{(l)}-\eta\nabla_{\theta}\mathcal{L}(\theta^{(l)},\bm{w}),$
where $\theta^{(l)}$ is the value of $\theta$ in $l$-th epoch and $\eta$ is the learning rate.
We now introduce our choices for the loss function and the termination criterion used in the above process.
Before introducing the loss function we used, we first introduce some properties of the eigenvalue problem (5) to give us an idea on how to choose the loss function.
Denote $V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}$ an arbitrary $k$-dimensional subspace of $V$ and $U_{k}=\text{span}\{u_{1},\cdots,u_{k}\}$ the eigensubspace corresponding to the first $k$ eigenvalues.
According to the the minimum-maximum principle [6], the sum of the first $k$ eigenvalues and eigenfunctions corresponding to the first $k$ eigenvalues satisfy the following equations,
@@ 3.3 Updating the basis functions
Motivated by (7) and (8), we then construct the following loss function for finding first $k$ eigenpairs (see also in [45]):
$\mathcal{L}(\theta,\bm{w})=\text{Trace}(\mathcal{B}^{-1}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,w_{k}))\mathcal{A}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,w_{k}))),$
where $v_{i}(x;\theta,w_{i})=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta)$ is an approximation of the eigenvector corresponding to the $k$-th eigenvalue, $i=1,\cdots,k.$
@@
As for the choice of termination criteria, we notice that the goal of training is to obtain improved basis functions compared to the initial ones, there is no need to train until the loss is small enough, e.g. find the $\theta^{\ast}$ s.t. $\mathcal{L}(\theta^{\ast},\bm{w})-\min\limits_{\theta\in\Theta}\mathcal{L}(\theta,\bm{w})\leq\epsilon$. Instead, in [49], the following termination criterion is adopted:
$\frac{\mathcal{L}(\theta,\bm{w})}{\mathcal{L}_{0}(\theta,\bm{w})}\leq\epsilon\ \text{or}\ N_{epoch}\geq N_{max},$
@@
where $\mathcal{L}_{0}(\theta,\bm{w})$ is the initial loss, $\epsilon$ is the tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum number of training epochs, respectively. However, since the minimum value of the loss function we used in this paper is not 0, this criterion may not be suitable for determining when to stop training. As an alternative, we propose a new termination criterion.
@@
We note that for an iterative method, the error between two adjacent iterations is usually used to determine when to terminate the iteration. That is, the iteration is terminated when the error between two adjacent iterations is less than some given tolerance $\epsilon$. Note that the loss usually decreases in an oscillatory manner during the training, hence using the change of the loss function during multiple adjacent epochs as an indicator may be better than using the change during two adjacent epochs. Therefore, we choose to use the change of the loss during multiple adjacent epochs as termination criterion. In this paper, we adopt the change of the loss during 10 adjacent epochs as an indicator. Any other reasonable choices are also acceptable.
@@
Let $l_{s}=\mathcal{L}_{s}(\theta,\bm{w})$ be the loss in the $s$-th epoch, $\bar{l}_{s}=\sum\limits_{i=1}^{10}|l_{s-i}$ $-l_{s+1-i}|$ be the change of the loss function during 10 adjacent epochs. The the following termination criterion is adopted:
@@
where $\epsilon$ is a given tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum training epochs, respectively. Under this criterion, the termination of training means that the relative change of the loss function during 10 adjacent epochs has been relatively small, which indicates that the benefits of continued training may be small. Since there is no need to train until the loss is small enough, a relatively big $\epsilon$ and a relatively small $N_{max}$ can be chosen.
3.4 Galerkin projection of the original eigenvalue problem onto the subspace
In this subsection, we will introduce the details on how to obtain the finite dimensional approximation of the eigenvalue problem (3) in the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$. From (3) in Section 2, we know that finding the Galerkin approximation of the eigenvalue problem (5) in subspace $,\text{span}\{\varphi_{1},\cdots$$,\varphi_{M}\}$ can be formulated as the following matrix eigenvalue problem:
$Aw=\lambda Bw,$ (10)
where $A=(a(\varphi_{i},\varphi_{j}))_{ij},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},w=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$

### v2

@@ 3 Subspace method based on neural networks for eigenvalue problems
$\begin{cases}Lu=\lambda u&\text{in }\rm\Omega,\\ u=0&\text{on }\partial\rm\Omega,\end{cases}$
where the differential operator $L:=-\nabla\cdot(\alpha\nabla)+\beta$ and $\alpha=(\alpha_{ij})_{ij}$ is symmetric positive definite with $\alpha_{ij}\in W^{1,\infty}({\rm\Omega})(i,j=1,\cdots,d)$, and $0\leq\beta\in L^{\infty}(\rm\Omega).$
The associated weak form of the eigenvalue problem above is as follows: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\rm\Omega)$ such that $b(u,u)=1$ and
$a(u,v)=\lambda b(u,v),\quad\forall v\in H_{0}^{1}(\rm\Omega),$ (5)
which is nothing but (1) with $a(u,v)=(\alpha\nabla u,\nabla v)+(\beta u,v)$ and $b(u,v)=(u,v)$. Here, $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ are two positive definite symmetric bilinear forms on $H_{0}^{1}({\rm\Omega})\times H_{0}^{1}(\rm\Omega)$.
@@
$\|v\|_{1}^{2}-C\|v\|_{0}^{2}\lesssim a(v,v),\quad\forall v\in H_{0}^{1}(\rm\Omega)$
@@
Let $V_{h}$ be a finite dimensional subspace of $V$. The finite dimensional discretization of (5) is as follows: Find $(\lambda,u_{h})\in\mathbb{R}\times V_{h}$ such that $b(u_{h},u_{h})=1$ and
$a(u_{h},v_{h})=\lambda b(u_{h},v_{h}),\quad\forall v_{h}\in V_{h}.$ (6)
In this paper, we consider to obtain the smallest $k$ eigenvalues and their corresponding eigenfunctions of the eigenvalue problem (5) and discretize the eigenvalue problem (5) in a finite dimensional space spanned by a neural network-based orthogonal basis. Our basic idea is as follows: we first use deep learning method to train some neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$ which is similar to [57]; then, we introduce some dimensionality reduction technique to obtain some orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ from these trained neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$; after that, we use some classical method to obtain the approximation solution of the eigenvalue problem (5) in the subspace $V_{h}=\text{span}\{\psi_{1},\cdots,\psi_{K}\}$. Usually, there holds $K\ll M$.
Here, we would like to point out that different from the work in [57], here, we add a step using some dimensionality reduction technique to the trained neural network-based functions. In fact, we have done some numerical tests, which show the necessity of using dimensionality reduction technique. Some of the results are provided in Appendix B.
We first give a general framework of our subspace method based on neural networks for solving eigenvalue problems, as in Algorithm 2, and then provide more details for each of the steps in the following parts of this section. Here, to make the framework more readable, we combine the training of the neural network-based functions using some deep learning method and obtaining the orthogonal basis from those trained function by some dimensionality reduction technique together and view them as a separate step for obtaining an orthogonal basis.
@@
Algorithm 2 General framework of subspace method based on neural networks for eigenvalue problems
 1. Construct a neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ by some deep learning method and dimensionality reduction technique.
 2. Use some classical method, e.g., the Galerkin method, to discretize the eigenvalue problem (5) in the subspace spanned by the basis $\{\psi_{i}\}_{i=1}^{K}$ and obtain an approximate solution.
3.1 Constructing a neural network-based orthogonal basis
In this subsection, we introduce how we obtain the neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ by using some deep learning method and dimensionality reduction technique.
Inspired by the subspace method based on neural networks proposed in [57], we consider to use neural network-based functions to construct basis and train the parameters of those neural network-based functions by minimizing some loss function. Furthermore, we observe the difference between solving eigenvalue problems and source problems. To solve the algebraic eigenvalue problem resulted from the operator eigenvalue problems with high accuracy, the stiffness matrix and the mass matrix generated by neural network-based basis functions are required to be symmetric and have good condition numbers. Usually, the trained neural network-based functions lack good linear independence. Consequently, some strategy must be taken to remove the redundant information contained in the trained neural network-based functions. In fact, as referred to in the beginning of this section, we have done some tests to show the necessity of using dimensionality reduction technique in Appendix B. In addition, to deal with boundary conditions, we also hope the basis satisfies boundary conditions of eigenvalue problem (5) exactly.
Based on the consideration above, we get the main steps for obtaining a neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ by using some deep learning method and dimensionality reduction technique as follows: First, choose a neural networks architecture with its parameters to be determined; then, modify neural network-based functions to make them satisfy boundary conditions exactly; after that, construct a proper loss function by the neural network-based functions and the information provided by the problem to be solved, and obtain the undetermined parameters in the neural network-based functions by minimizing the loss function; finally, construct a neural network-based orthogonal basis from the trained neural network-based functions by using some dimensionality reduction technique. In the following of this subsection, we will introduce the details for each step.
3.1.1 Neural networks architecture
The neural networks architecture used here is similar to the one introduced in the previous section. The only difference is that the neural networks in Section 2 is used to approximate the solution to the source problem, here the neural networks are used to approximate the eigenfunctions corresponding to the first $k$ eigenvalues of some partial differential operator. Therefore, the dimension of the output layer is $k$. Thus, the output of the neural networks can be presented as follows,
$\bm{u}(\bm{x};\theta,w)=(\bm{\varphi}\cdot\bm{w}_{1},\cdots,\bm{\varphi}\cdot\bm{w}_{k})^{T}=\left(\sum\limits_{j=1}^{M}w_{1}^{(j)}\varphi_{j}(\bm{x},\theta),\cdots,\sum\limits_{j=1}^{M}w_{k}^{(j)}\varphi_{j}(\bm{x},\theta)\right)^{T},$
where $w=(\bm{w}_{1},\cdots,\bm{w}_{k})$, $\bm{w}_{s}=(w_{s}^{(1)},\cdots,w_{s}^{(M)})^{T}$ and $u_{s}(\bm{x};\theta,\bm{w}_{s})$ denotes the $s$-th output of the neural networks for $s=1,\cdots,k$. Similar to those introduced in Section 2, here, $\bm{\varphi}(\bm{x},\theta)=\sigma\circ\mathcal{T}_{L+1}\circ\cdots\circ\sigma\circ\mathcal{T}_{1}(\bm{x})$, $\sigma$ is an activation function, $\mathcal{T}_{l}(\bm{y})=W_{l}\cdot\bm{y}+\bm{b}_{l}$ and $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}$ are the parameters of the neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$.
@@ 3.1.1 Neural networks architecture
The initialization of the parameters $\theta$ and $\bm{w}_{1},\cdots,\bm{w}_{k}$ are as follows. For the parameters $\theta$, we adopt the default initialization strategy provided by PyTorch. For the parameters $\bm{w}_{1},\cdots,\bm{w}_{k}$, there are various methods for assigning initial values. Numerical experiments provided in [57] indicate that different choices of $\bm{w}_{1},\cdots,\bm{w}_{k}$ have no obvious impact on the results. Therefore, here, we do not pay much attention on the choice of $\bm{w}_{1},\cdots,\bm{w}_{k}$. We just give some simple choices. That is, for a single output, we choose $\bm{w}=(1,\cdots,1)^{T}\in\mathbb{R}^{M}$. For multiple outputs, a random approach is chosen, that is, $\bm{w}_{1},\cdots,\bm{w}_{k}\in\mathbb{R}^{M}$ are assigned random values generated from a uniform distribution defined over the interval [1,1][-1,1].
3.1.2 Handling the boundary condition
@@
For eigenvalue problems, some boundary conditions are usually required. Since the neural network-based functions are usually nonlocal, it is not so easy to deal with the boundary condition by using the same way as classical methods. Here, we choose to modify those functions to make them satisfy the boundary condition. For example, for the case of zero boundary condition with $\rm\Omega$ being bounded, we can modify the basis function by the following way. We multiply $\varphi_{i}$ by a function $f$ such that
$\bar{\varphi}_{i}=\varphi_{i}f,\ \bar{\varphi}_{i}(\bm{x})=0,\ \forall\bm{x}\in\partial{\rm\Omega},\ \forall i=1,\cdots,M.$ (7)
@@
If $\rm\Omega$ is unbounded, it is necessary to modify the neural network-based functions to ensure that the modified functions are $L^{2}$ integrable. We can achieve this by multiplying $\varphi_{i}$ by a function $f$ such that
$\bar{\varphi}_{i}=\varphi_{i}f,\ \int_{\rm\Omega}\bar{\varphi}_{i}^{2}dx<\infty,\ \forall i=1,\cdots,M.$ (8)
Since the Hermite-Gauss quadrature scheme is used to numerically approximate the integrals on an unbounded domain in this paper, we can choose $f(\bm{x})=\exp(-\frac{1}{2}\bm{x}^{T}\bm{x})$.
For simplicity, in the remaining parts of this paper, $\varphi_{i}$ will refer to the modified neural network-based function $\bar{\varphi}_{i}$ as described above.
3.1.3 The loss function and the training
In this subsection, we will introduce our choice for the loss function $\mathcal{L}$ and a specific approach to train the neural network-based functions.
We first introduce our choices for the loss function. Before that, we first introduce some properties of the eigenvalue problem (5) to give us an idea on how to choose the loss function.
Denote $V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}$ an arbitrary $k$-dimensional subspace of $V$ and $U_{k}=\text{span}\{u_{1},\cdots,u_{k}\}$ the eigensubspace corresponding to the smallest $k$ eigenvalues.
According to the minimum-maximum principle [6], the sum of the smallest $k$ eigenvalues and eigenfunctions corresponding to the smallest $k$ eigenvalues satisfy the following equations,
@@ 3.1.3 The loss function and the training
We now consider the approximation of the eigenvalue problem (5) on the finite dimensional subspace $V_{h}$ spanned by neural network-based functions. Set $V_{h}=\text{span}\{\varphi_{1}(\bm{x};\theta),\cdots,\varphi_{M}(\bm{x};\theta)\}\subset H^{1}_{0}(\rm\Omega)$. Set $v_{i}(\bm{x};\theta,\bm{w}_{i})=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(\bm{x};\theta)$, $i=1,\cdots,k$. Motivated by (9) and (10), we then construct the following loss function for finding first $k$ eigenpairs (see also in [53]):
$\mathcal{L}(\theta,w)=\text{Trace}(\mathcal{B}^{-1}(v_{1}(x;\theta,\bm{w}_{1}),\cdots,v_{k}(x;\theta,\bm{w}_{k}))\mathcal{A}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,\bm{w}_{k}))).$ (11)
Now, we turn to introduce how we train the neural network-based functions. Similar to Algorithm 1 proposed in [57], for training the neural network-based functions, we first fix the parameter $w$, that is, set $w=w^{(0)}$ with $w^{(0)}$ being given at initialization and then obtain parameters $\theta$ by solving the following minimization problem:
$\min_{\theta\in\rm\Theta}\mathcal{L}(\theta,w^{(0)}).$ (12)
The gradient descent method is adopted to solve the minimization problem (12). That is, we update the parameters $\theta$ by
$\theta^{(l+1)}=\theta^{(l)}-\eta\nabla_{\theta}\mathcal{L}(\theta^{(l)},w^{(0)}),$
where $\theta^{(l)}$ is the value of $\theta$ in $l$-th epoch and $\eta$ is the learning rate adjusted by the Adam optimizer [28]. Here, when minimizing the loss function, a fixed $w^{(0)}$ is adopted
We have also briefly explained how the training process affects the accuracy of final solution through some numerical experiments in Appendix A.
@@
For the choice of termination criteria of the gradient descent method, we notice that the goal of training is to obtain the basis other than the solution, there is no need to train until the loss is small enough, e.g. find the $\theta^{\ast}$ s.t. $\mathcal{L}(\theta^{\ast},w^{(0)})-\min\limits_{\theta\in\rm\Theta}\mathcal{L}(\theta,w^{(0)})\leq\epsilon$. Instead, in [57], the following termination criterion is adopted:
$\frac{\mathcal{L}(\theta,w^{(0)})}{\mathcal{L}_{0}(\theta,w^{(0)})}\leq\epsilon\ \text{or}\ N_{epoch}\geq N_{max},$
@@
where $\mathcal{L}_{0}(\theta,w^{(0)})$ is the initial loss, $\epsilon$ is the tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum number of training epochs, respectively. However, since the minimum value of the loss function we used in this paper is not 0, this criterion may not be suitable for determining when to stop training. As an alternative, we propose a new termination criterion.
@@
We note that for an iterative method, the error between two adjacent iterations is usually used to determine when to terminate the iteration. That is, the iteration is terminated when the error between two adjacent iterations is less than some given tolerance $\epsilon$. Note that the loss usually decreases in an oscillatory manner during the training, hence using the change of the loss function during multiple adjacent epochs as an indicator may be better than using the change during two adjacent epochs. Therefore, we choose to use the change of the loss during multiple adjacent epochs as termination criterion. In this paper, we adopt the change of the loss during $m$ adjacent epochs as an indicator. Any other reasonable choices are also acceptable.
@@
Let $l_{s}=\mathcal{L}_{s}(\theta,w)$ be the loss in the $s$-th epoch, $\bar{l}_{s}=\sum\limits_{i=1}^{m}|l_{s-i}$ $l_{s+1-i}|$ be the change of the loss function during $m$ adjacent epochs. The following termination criterion is adopted:
@@
where $\epsilon$ is a given tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum training epochs, respectively. Under this criterion, the termination of training means that the relative change of the loss function during $m$ adjacent epochs has been relatively small, which indicates that the benefits of continued training may be small. Since there is no need to train until the loss is small enough, a relatively big $\epsilon$ and a relatively small $N_{max}$ can be chosen.
3.1.4 Dimensionality reduction technique
In this subsection, we will introduce how we obtain the neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ from the trained neural networks-based functions $\{\varphi_{i}\}_{i=1}^{M}$ which may numerically lack good linear independence while keeping as much necessary information as possible by some dimensionality reduction technique. Here, we choose the proper orthogonality decomposition (POD) to achieve this goal.
We can describe the problem as follows. Denote $V=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, $V_{k}\subset V$ any $K$-dimensional subspace of $V$, and ${\rm\Pi}_{K}$ the orthogonal projection operator from $V$ to $V_{K}$. The problem is to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate the original space $V$ good enough. There will be two subproblems. One is that, for a given dimension of the subspace $K\leq M$, how to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate the original space $V$ in a least–squares sense, the other is how to choose a suitable dimension of the subspace $K$. For the first problem, i.e. we seek an optimal orthogonal projection ${\rm\Pi}_{K}^{\ast}:V\rightarrow V_{K}^{\ast}$ minimizing the projection error of ${\rm\Phi}=(\varphi_{1},\cdots,\varphi_{M})^{T}\in[L^{2}({\rm\Omega})]^{M}$ under the following least–squares distance,
[…]
Hence, we need to solve the following minimization problem,
[…]
Denote $U=[u_{1},\cdots,u_{K}]\in\mathbb{R}^{M\times K}$ and
${\rm\Psi}=U^{T}{\rm\Phi}.$ (15)
We can rewrite the minimization problem (14) to the minimization problem (16) through the following simple derivation,
[…]
Notice that the problem (16) is nothing but the minimization problem considered in the POD method, which has been systematically studied (see [39, 50]), hence we give the following optimal orthogonal projection ${\rm\Pi}_{K}^{\ast}$ of the problem (16) directly,
$\mathcal{M}^{2}u_{j}=\mu_{j}\mathcal{M}u_{j},j=1,\cdots,K,$ (17)
[…]
where $\mu_{1}\geq\cdots\mu_{M}\geq 0$ are the first $M$ eigenvalues of $\mathcal{M}^{2}$ relative to $\mathcal{M}$, $u_{1},\cdots,u_{M}$ are the corresponding eigenvectors.
$\mathcal{M}\bar{u}_{j}=\mu_{j}\bar{u}_{j},j=1,\cdots,K,$
$u_{j}=\frac{1}{\sqrt{\mu_{j}}}\bar{u}_{j},j=1,\cdots,K.$
Remark 2
In actual implementation, the condition number of $\mathcal{M}$ is usually very large, and it is difficult to obtain a high-precision solution by directly solving the eigenvalue problem. Instead, we choose to calculate the singular value of $\mathcal{M}^{\frac{1}{2}}$, from which we can also get $\mu_{j},u_{j},j=1,\cdots,K.$ By the way, we only need to deal with a problem whose condition number becomes the square root of the original problem.
In addition, we do not need to pay any extra cost for calculating $\mathcal{M}^{\frac{1}{2}}$, because it has a natural approximation as follows. Denote $\Phi_{h}=(\varphi_{i}(\bm{x}^{(j)}))_{ij}\in\mathbb{R}^{M\times n},W_{h}=diag(\rho_{1},\cdots,\rho_{n})\in\mathbb{R}^{n\times n}$, where $\{\bm{x}^{(i)},\rho_{i}\}_{i=1}^{n}$ represents the numerical integration point and weight, respectively. For $\mathcal{M}$, we have the following approximation, $\mathcal{M}\approx\Phi_{h}W_{h}\Phi_{h}^{T}$. Hence, we can use $W_{h}^{\frac{1}{2}}\Phi_{h}^{T}$ to approximate $\mathcal{M}^{\frac{1}{2}}$.
Regarding the choice of the subspace dimension $K$, since (18) shows the approximate error of the optimal $K$-dimensional subspace is related to the sum of the last MKM-K eigenvalues of the matrix $\mathcal{M}$ (or the square root of the eigenvalues), it’s reasonable to choose a small enough $K$ such that the first $K$ eigenvalues (or the square root of the eigenvalues) account for a high proportion in the overall, i.e.,
$K=\mathop{\arg\min}\{I(N):I(N)\geq 1-\gamma\},$
where $I(N)=\frac{\sum\limits_{j=1}^{N}\mu_{j}}{\sum\limits_{j=1}^{M}\mu_{j}}$ or $I(N)=\frac{\sum\limits_{j=1}^{N}\sqrt{\mu_{j}}}{\sum\limits_{j=1}^{M}\sqrt{\mu_{j}}}$ and $\gamma$ is a given parameter.
In our experiment, we find that the indicator $I(N)$ defined by the latter works better.
To see the necessity of employing dimensionality reduction technique, we have done some numerical tests, whose results are present in Appendix B.
3.1.5 Algorithm
By summarizing the steps above, we obtain our algorithm for constructing the neural network-based orthogonal basis by deep learning method and dimensionality reduction technique, see Algorithm 3 for the details.
Algorithm 3 Constructing the neural network-based orthogonal basis
0: the neural network structure: $L,n_{0},n_{1},\cdots,n_{L},M$; parameters used for training the basis: $\epsilon,m,N_{\max},\gamma$; number of eigenpairs needed: $k$
0:  the number of orthogonal basis functions $K$ and the basis $\{\psi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{K}.$
 1. Choose a neural network architecture and modify neural network-based functions to make it satisfy the boundary condition.
 2. Train the neural network-based functions $\{\varphi_{i}(\bm{x},\theta)\}_{i=1}^{M}$ by minimizing the loss function (11) until (13) is satisfied.
 3. Obtain the neural network-based orthogonal basis $\{\psi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{K}$ from the trained neural network-based functions $\{\varphi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{M}$ by POD method.
3.2 Galerkin discretization in the neural network-based subspace
In this subsection, we will introduce the details on how we discretize the eigenvalue problem (5) in the subspace spanned by the neural network-based orthogonal basis $\{\psi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{K}$.
Similar to the introduction in Section 2, we consider the following finite dimensional discretization of (5): Find $(\lambda,u_{h})\in\mathbb{R}\times V_{h}$ such that $b(u_{h},u_{h})=1$ and
$a(u_{h},v_{h})=\lambda b(u_{h},v_{h}),\quad\forall v_{h}\in V_{h}.$ (19)
Let $V_{h}=\text{span}\{\psi_{1}(\bm{x};\theta^{\ast}),\cdots,\psi_{K}(\bm{x};\theta^{\ast})\}$. Set $u_{NN}(\bm{x};\theta^{\ast},\bm{w})=\sum\limits_{j=1}^{K}w^{(j)}\psi_{j}(\bm{x};\theta^{\ast})$, and substitute $u_{NN}(\bm{x};\theta^{\ast},\bm{w})$ into (19) and choose $v_{h}(\bm{x})=\psi_{i}(\bm{x};\theta^{\ast})$, we get
$\sum\limits_{j=1}^{K}w^{(j)}a(\psi_{j},\psi_{i})=\lambda\sum\limits_{j=1}^{K}w^{(j)}b(\psi_{j},\psi_{i}),i=1,\cdots,K.$ (20)
If we set $A=(a(\psi_{i}(\bm{x};\theta^{\ast}),\psi_{j}(\bm{x};\theta^{\ast})))_{ij},B=((\psi_{i}(\bm{x};\theta^{\ast}),\psi_{j}(\bm{x};\theta^{\ast})))_{ij}\in\mathbb{R}^{K\times K},$ $\bm{w}=(w^{(1)},\cdots,w^{(K)})^{T}\in\mathbb{R}^{K}$, then (20) can be written as the following algebraic eigenvalue problem
$A\bm{w}=\lambda B\bm{w}.$ (21)